OpenStudy (anonymous):

hey i need help The sum of the squares of 2 consecutive negative integers is 41. What are the numbers? Which of the following equations is the result of using the factoring method to solve the problem? (n - 5)(n - 4) = 0 (n - 5)(n + 4) = 0 (n + 5)(n - 4) = 0 (n + 5)(n + 4) = 0

OpenStudy (raden):

let the both numbers are -n and -(n+1)

OpenStudy (raden):

so, (-n)^2 + (-(n+1))^2 = 41 n^2 + (n+1)^2 = 41 n^2 + n^2 + 2n + 1 - 41 = 0 2n^2 + 2n - 40 = 0 both sides just divide by 2, n^2 + n - 20 = 0 (n + 5)(n - 4) = 0

OpenStudy (anonymous):

now can you help me with this one The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem? 3n2+ 5 = 116 3n2+ 3n + 3 = 116 3n2+ 6n + 5 = 116

OpenStudy (raden):

let that numbers are : n, n+1 and n+2. so, n^2 + (n+1)^2 + (n+2)^2 = 116 n^2 + n^2 + 2n + 1 + n^2 + 4n + 4 = 116 3n^2 + 6n + 5 = 116

OpenStudy (anonymous):

i dont have an answer like that

OpenStudy (raden):

the last option, right ?

OpenStudy (anonymous):

never mind its right

OpenStudy (anonymous):

ok can you help me with some more

OpenStudy (raden):

but, i think there is a problem with that question...

OpenStudy (anonymous):

no it says its right

OpenStudy (anonymous):

The square of a number is equal to 10 less than 7 times that number. What are the two possible solutions? Which of the following equations is used in the process of solving this problem? x2- 7x + 70 = 0 x2- 7x + 10 = 0 x2+ 7x - 10 = 0 i need help with this one

OpenStudy (raden):

if i want to solve for n, 3n^2 + 6n + 5 = 116 3n^2 + 6n + 5 - 116 = 0 3n^2 + 6n - 111 = 0 both sides divide by 3, n^2 + 2n - 37 = 0 it can not be factorizing, it means that the numbers are not integer number!!

OpenStudy (raden):

so, the 2nd is a wrong question (statemen)

OpenStudy (raden):

*statement

OpenStudy (anonymous):

ok iam lost now

OpenStudy (raden):

n^2 + 2n - 37 = 0 to find the its solution, we use quadratic formula :